How many liters of 24% salt solution must be mixed with 24 liters of a 40% salt solution to get a 34% salt solution?
Answer by mananth(16946)

-------------- percent ---------------- Amount
Salt solution I 40 ---------------- 24 liters
Salt solution II 24 ---------------- x liters
Mixture 34 ---------------- 24 + x liters

40*24+24x=34(24 +x)
960+24x =816+34x
24x-34x =-960+ 816
-10x=-144
/ -10
x = 14.4 liters Salt solution II

m.ananth@hotmail.ca